40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3741     Entry time: Tue Oct 19 15:14:51 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: PRM (little) update 

[Jenne, Suresh]

We've aligned the guiderod and wire standoff to the PRM, each partly.  They have both been aligned to the correct distance above the scribe lines, but they have not yet been centered forward/backward along the thickness of the optic.  So, we're working on it...

ELOG V3.1.3-